Orthogonal
Is it true that if x is orthogonal to v and w, then x is also orthogonal to v − w. Why so?
1
Expert's answer
2020-11-11T08:52:09-0500

It is well known that the scalar product is a distributive with regard to the sum and the difference of vectors:


"a\cdot(b\pm c)=a\cdot b\pm a\cdot c".


The vectors "a" and "b" are ortogonal if and only if the scalar product "a\cdot b" is equal to 0.


If  "x" is orthogonal to "v" and "w", then "x\cdot v=0" and "x\cdot w=0". Therefore,


"x\cdot(v-w)=x\cdot v- x\cdot w =0-0=0,"


and we conclude that "x" is also orthogonal to "v \u2212 w."